# Policy Optimization with Smooth Guidance Learned from State-Only Demonstrations

### v2

@@ Abstract
The sparsity of reward feedback remains a challenging problem in online deep reinforcement learning (DRL). Previous approaches have utilized offline demonstrations to achieve impressive results in multiple hard tasks. However, these approaches place high demands on demonstration quality, and obtaining expert-like actions is often costly and unrealistic. To tackle these problems, we propose a simple and efficient algorithm called Policy Optimization with Smooth Guidance (POSG), which leverages a small set of state-only demonstrations (where only state information is included in demonstrations) to indirectly make approximate and feasible long-term credit assignments and facilitate exploration. Specifically, we first design a trajectory-importance evaluation mechanism to determine the quality of the current trajectory against demonstrations. Then, we introduce a guidance reward computation technology based on trajectory importance to measure the impact of each state-action pair. We theoretically analyze the performance improvement caused by smooth guidance rewards and derive a new worst-case lower bound on the performance improvement. Extensive results demonstrate POSG’s significant advantages in control performance and convergence speed in four sparse-reward environments, including the grid-world maze, Hopper-v4, HalfCheetah-v4, and Ant maze. Notably, the specific metrics and quantifiable results are investigated to demonstrate the superiority of POSG.
keywords:
@@
1 Introduction
In recent years, deep reinforcement learning (RL) has demonstrated remarkable accomplishments in tackling sequential decision-making challenges across diverse domains, including the Arcade Learning Environment Mnih2015HumanlevelCT ; bellemare2013arcade ; ecoffet2021first , the game of Go Silver2016MasteringTG , continuous locomotive control Lillicrap2016ContinuousCW ; Schulman2015TrustRP ; fujimoto2018addressing , and robotic navigation Florensa2017StochasticNN ; zhanghierarchical ; li2022active . Despite these celebrated achievements, reinforcement learning remains a formidable task in scenarios characterized by sparse or delayed rewards yang2021exploration ; gangwani2020learning , primarily due to the challenge of striking a balance between exploration and exploitation in environments with sparse or delayed rewards guo2020memory . Conventional deep RL algorithms confront exploration difficulties, as fully exploring the entire state-action space is often unfeasible and cannot be spontaneously guaranteed, particularly in settings with sparse environmental rewards jing2020reinforcement .
One possible solution to the problem of sparse or delayed rewards is temporal credit assignment. Credit assignment aims to understand the relevance between actions and outcomes and measure the impact of actions the agent performs on future rewards. Some recent studies propose building an environmental model to obtain a more fine-grained description of the action’s effect Ha2018WorldM ; Kaiser2019ModelBasedRL ; Schrittwieser2019MasteringAG . However, this approach increases the computational burden, and acquiring an accurate model in complex and partially observed environments remains difficult. On the other hand, model-free credit assignment methods leverage hindsight information harutyunyan2019hindsight , counterfactual mesnard2021counterfactual , episode memory Hung2018OptimizingAB , transformers Ferret2019CreditAA , and return decomposition arjona2019rudder to perform long-term credit assignments. However, the premise of these methods is to obtain good trajectories with sparse rewards, and these designated methods might cause the problem of unstable training and parameter sensibility.
Many research advances have assisted policy exploration of agents by learning from demonstrations (LfD) zhou2020watch . An intuitive LfD approach enhances RL by data augmentations, which maintains the expert demonstrations in a replay buffer for value estimation hester2018deep ; vecerik2017leveraging ; gulcehre2019making ; libardi2021guided . Some LfD methods utilize demonstrations to pre-train the policy by supervised learning silver2016mastering ; cruz2017pre . These algorithms force the agent’s policy to conform to the expert’s policy and do not reuse them during the policy optimization procedure. Recent LfD studies draw inspiration from imitation learning and encourage the agent to mimic the demonstrated actions kang2018policy ; Sun2018TruncatedHP ; jing2020reinforcement ; rengarajan2022reinforcement . Specifically, these methods either augment the original RL loss function with a divergence regularization term or design a new shaping reward derived from a distribution divergence function to force expert-alike exploration.
@@ 1 Introduction
The proposed approach leverages only a few, or even a single state-only demonstration trajectory, to indirectly estimate the impact of each state-action pair. The fundamental insight lies in utilizing state-only demonstration trajectories to assess the impact of current state-action pairs, thereby amalgamating state distribution information from demonstrations with their associated reward signals. We introduce two technologies to compute the impact of each state-action pair: a trajectory-importance evaluation mechanism and a guidance reward computation technology. Specifically, the trajectory-importance evaluation mechanism estimates trajectory importance based on the maximum mean discrepancy (MMD) distance to demonstrations and the returns of corresponding trajectories. Subsequently, the guidance reward for each state-action pair is derived through a smooth weighted average of trajectory importance. Furthermore, we establish a novel worst-case lower bound for POSG policy optimization and offer a theoretical guarantee of performance improvement. Extensive experimental evaluations demonstrate that POSG surpasses other baseline algorithms across a discrete grid-world maze and three continuous locomotion control tasks.
@@
    POSG can use a single state-only demonstration trajectory to solve RL’s exploration problem and improve RL’s sample efficiency, demonstrated experimentally in Section 7.4.1.
@@
The rest of this article is organized as follows. Section 2 describes some important progress of the recent related work. Section 3 briefly introduces the related preliminary knowledge. Then, the method we propose is introduced in detail in Section 4. Section 5 describes the theoretical analyses of POSG. The environmental settings are introduced in Section 6. In Section 7, we experimentally demonstrate the feasibility and effectiveness of the proposed algorithm in terms of exploration efficiency and learning speed. Finally, we summarize the main work of this study in Section 8.
2 Related Work
@@
Learning from Demonstrations. LfD combines RL with expert demonstration data to improve policy exploration and accelerate learning. LfD trains a behavioral policy with a state-action visitation distribution similar to the demonstrator’s to accomplish this goal. Early LfD algorithms sample replay buffers containing expert demonstrations and self-generated data simultaneously to enhance the learning ability of the agent hester2018deep ; vecerik2017leveraging . Self-imitation learning (SIL) methods oh2018self ; gangwani2019learning ; libardi2021guided train the agent to imitate its own past experiences only when the return of the previous episode is greater than the value estimate of the agent or returns of trajectories in the replay buffer. Episodic reinforcement learning methods utilize episodic memories to estimate the value of the state precisely and propagate its value to the previous states pritzel2017neural ; blundell2016model ; lin2018episodic ; lee2019sample ; hu2021generalizable ; le2021model . Many other previous works introduce novel methods that enable the agent learns a range of diverse exploratory policies based on episodic memory guo2020memory ; badia2020never . This study uses demonstration experiences in the state-action space to construct a shaping function, which can easily be integrated with existing RL methods, such as PPO schulman2017proximal .
@@
State-Only Imitation Learning (IL). Imitation learning aims to learn a control policy that imitates the behaviors of experts and outputs the same action when the agent receives the observation that occurred in the demonstration data set. The state-only IL method is a branch of imitation learning, where the requirement for the demonstrator’s action information is alleviated. This approach expands the scope of the realistic application of IL. GAIfO torabi2018generative proposes to learn a policy to output actions that lead to similar effects as demonstrations. I2L Gangwani2020StateonlyIW and SAIL Liu2019StateAI train the agent by minimizing the distance between state visitation distributions of the current policy and demonstrations. AILO Gangwani2022ImitationLF leverages demonstrations only containing observations to train an intermediary policy whose state transitions are close to the expert dataset.
@@
Metrics to Compute the Distance between Policies or Trajectories. Many studies propose to compute the distance between policies using Kullback–Leibler (KL), Bregman, or $f$-divergence lee2019efficient ; ghasemipour2020divergence ; wang2019divergence . Rényi divergence is introduced to compute the discrepancy of state visitation between different trajectories Yuan2022RewardingEV ; however, this divergence is a parametric distance measure, and this approach employs $k$-NN estimator to efficiently estimate the parameters of distributions. MADE maximizes the derivation of state-action visitations of the current policy from the explored regions of previous trajectories zhang2021made . However, MADE relies on the estimation of the state-action visitation density, which can be non-trivial in high-dimensional control tasks. Similar to our method, CQL kumar2020conservative and MCPO jing2020reinforcement adopt the MMD metric as the discrepancy measure; however, these algorithms require both state and action information of demonstrations to update the policy parameter.
@@
Credit Assignment. Various works have focused on the credit assignment problem, and credit assignment methods can be integrated with existing RL algorithms easily pocius2018comparing . These methods assist us in understanding the association between sparse or delayed rewards and state or state-action pairs and reduce learning time by providing dense and supplemental rewards ng1999policy . Zheng et al. zheng2022adaptive formulate pairwise weight functions of the state where the action is performed, the future highly-rewarded state, and the time horizon between the two states, which is learned by a special meta-gradient. Hindsight Credit Assignment (HCA) harutyunyan2019hindsight converts the credit assignment problem into a supervised learning task by learning a hindsight probability function of actions. Counterfactual Credit Assignment (CCA) mesnard2021counterfactual uses a value function baseline with a hindsight information vector to implement credit assignment implicitly while avoiding giving away information about the agent’s actions to reduce potential bias. In RUDDER arjona2019rudder , an LSTM network is used to learn to redistribute the return of a trajectory to the preceding states before the final rewarding state. Episodic Backward Update lee2019sample and Neural Episodic Control pritzel2017neural enable efficient reward propagation by sampling from episodic memories and updating the value of all transitions more quickly. State Associative Learning raposo2021synthetic propagates credit directly by learning associations between states and arbitrary future states. Xu et al. xu2018meta introduce an online meta-learning method to learn hyper-parameters of a discount $\gamma$ and bootstrapping parameter $\lambda$ of Temporal Difference (TD) for credit assignment. In contrast, our method does not incur high computational costs by dispensing with training auxiliary networks and can be regarded as a simple weighted return decomposition.
3 Preliminaries
@@
This paper studies the credit assignment problem of reinforcement learning in tasks with sparse or delayed rewards. Before describing our method in detail, we introduce some preliminary knowledge about RL in this section. Then, the definition of Maximum Mean Discrepancy (MMD), which plays an important role in our method, is provided.
3.1 Reinforcement Learning
@@
3.2 Maximum Mean Discrepancy
In this paper, we treat trajectories as deterministic policy distributions and use the maximum mean discrepancy (MMD) to measure the difference (or similarity) between trajectories empirically gretton2006kernel ; Gretton2012OptimalKC ; masood2019diversity . MMD is a non-parametric test statistic, and different from Kullback-Leibler (KL), Jensen–Shannon (JS), or $f$-divergences, calculating the MMD metric does not require the distribution parameters. This property makes it suitable for the problem considered in this study where trajectories are treated as deterministic policies.
Given any two trajectories $\tau$ and $\upsilon$, we can use the maximum mean discrepancy to measure the distance between them:
$\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})=\underset{x,x^{\prime}\sim\rho_{\tau}}{\mathbb{E}}\left[k\left(x,x^{\prime}\right)\right]-2\underset{\begin{subarray}{c}x\sim\rho_{\tau}\\ y\sim\rho_{\tau_{E}}\end{subarray}}{\mathbb{E}}\left[k(x,y)\right]+\underset{y,y^{\prime}\sim\rho_{\tau_{E}}}{\mathbb{E}}\left[k(y,y^{\prime})\right],$ (5)
where $k(\cdot,\cdot)$ is the kernel of a reproducing kernel Hilbert space $\mathcal{H}$. In practice, the function $k(\cdot,\cdot)$ in Eq. (5) is often defined as:
$k(x,y)=K\left(g(x),g(y)\right).$ (6)
Formally, the function $g$ gives us the feasibility that adapts the focus of the ${\rm MMD}$ metric for different aspects, such as state visits, action choices, or both, to different downstream tasks. In our experiments, we define the distance $D(\tau,\mathcal{M}_{E})$ from trajectory to the demonstration dataset $\mathcal{M}_{E}$ as follows:
This definition of $D(\tau,\mathcal{M}_{E})$ only concerns a relevant subset of the information in state-action pairs. For example, we choose this information subset to be the coordinates $c=(x,y)$ of its center of mass (CoM) in the Key-Door-Treasure domain, i.e., the function $g$ maps a state-action pair $(s,a)$ to $c=(x,y)$.
4 Proposed Approach
This section presents Policy Optimization with Smooth Guidance (POSG), a computation-efficient framework for achieving approximate long-term credit assignments. The key insight is that we propose to exploit state-only demonstration trajectories to measure the relative importance of current state-action pairs. Our approach merges the distributional information of demonstrations with the return information of trajectories.
4.1 Smooth Guidance Reward Design
This section defines the trajectory importance based on a novel trajectory-level distance measurement and the trajectory return. In this manner, this importance integrates the state distribution information of demonstrations with reward signals of trajectories. We can then fuse these two pieces of information into the guidance reward by a smooth weighted average of the trajectory importance.
@@ 4.1 Smooth Guidance Reward Design
We then introduce a trajectory weight function $\omega_{\pi}(\tau|\mathcal{M}_{E})$ in Eq. (10) based on the MMD distance metric over the trajectory space and the conditional distribution $p_{\pi}(\tau|x)$. Mathematically, an exponential function $e^{-kd}$ is adopted by $\omega_{\pi}(\cdot|\mathcal{M}_{E})$ to smooth the trajectory weight, and $\omega_{\pi}(\cdot|\mathcal{M}_{E})$ is expressed as follows:

### v3

@@ Abstract
The sparsity of reward feedback remains a challenging problem in online deep reinforcement learning (DRL). Previous approaches have utilized offline demonstrations to achieve impressive results in multiple hard tasks. However, these approaches place high demands on demonstration quality, and obtaining expert-like actions is often costly and unrealistic. To tackle these problems, we propose a simple and efficient algorithm called Policy Optimization with Smooth Guidance (POSG), which leverages a small set of state-only demonstrations (where expert action information is not included in demonstrations) to indirectly make approximate and feasible long-term credit assignments and facilitate exploration. Specifically, we first design a trajectory-importance evaluation mechanism to determine the quality of the current trajectory against demonstrations. Then, we introduce a guidance reward computation technology based on trajectory importance to measure the impact of each state-action pair, fusing the demonstrator’s state distribution with reward information into the guidance reward. We theoretically analyze the performance improvement caused by smooth guidance rewards and derive a new worst-case lower bound on the performance improvement. Extensive results demonstrate POSG’s significant advantages in control performance and convergence speed in four sparse-reward environments, including the grid-world maze, Hopper-v4, HalfCheetah-v4, and Ant maze. Notably, the specific metrics and quantifiable results are investigated to demonstrate the superiority of POSG.
Index Terms:
@@
I Introduction
In recent years, deep reinforcement learning (RL) has demonstrated remarkable accomplishments in tackling sequential decision-making challenges across diverse domains, including the Arcade Learning Environment [1, 2, 3], the game of Go [4], continuous locomotive control [5, 6, 7], and robotic navigation [8, 9, 10]. Despite these celebrated achievements, reinforcement learning remains a formidable task in scenarios characterized by sparse or delayed rewards [11, 12], primarily due to the challenge of striking a balance between exploration and exploitation in environments with sparse or delayed rewards [13]. Conventional deep RL algorithms confront exploration difficulties, as fully exploring the entire state-action space is often unfeasible and cannot be spontaneously guaranteed, particularly in settings with sparse environmental rewards [14].
One possible solution to the problem of sparse or delayed rewards is temporal credit assignment. Credit assignment aims to understand the relevance between actions and outcomes and measure the impact of actions the agent performs on future rewards. Some recent studies propose building an environmental model to obtain a more fine-grained description of the action’s effect [15, 16, 17]. However, this approach increases the computational burden, and acquiring an accurate model in complex and partially observed environments remains difficult. On the other hand, model-free credit assignment methods leverage hindsight information [18], counterfactual [19], episode memory [20], transformers [21], and return decomposition [22] to perform long-term credit assignments. However, the premise of these methods is to obtain good trajectories with sparse rewards, and these designated methods might cause the problem of unstable training and parameter sensibility.
Many research advances have assisted policy exploration of agents by learning from demonstrations (LfD) [23]. An intuitive LfD approach enhances RL by data augmentations, which maintains the expert demonstrations in a replay buffer for value estimation [24, 25, 26, 27]. Some LfD methods utilize demonstrations to pre-train the policy by supervised learning [28, 29]. These algorithms force the agent’s policy to conform to the expert’s policy and do not reuse them during the policy optimization procedure. Recent LfD studies draw inspiration from imitation learning and encourage the agent to mimic the demonstrated actions [30, 31, 14, 32]. Specifically, these methods either augment the original RL loss function with a divergence regularization term or design a new shaping reward derived from a distribution divergence function to force expert-alike exploration.
@@ I Introduction
The proposed approach leverages only a few, or even a single state-only demonstration trajectory, to indirectly estimate the impact of each state-action pair. The fundamental insight lies in utilizing state-only demonstration trajectories to assess the impact of current state-action pairs, thereby amalgamating state distribution information from demonstrations with their associated reward signals. We introduce two technologies to compute the impact of each state-action pair: a trajectory-importance evaluation mechanism and a smooth guidance reward computation technology. Specifically, the trajectory-importance evaluation mechanism estimates trajectory importance based on the maximum mean discrepancy (MMD) distance to demonstrations and the returns of corresponding trajectories. Subsequently, the guidance reward for each state-action pair is derived through a smooth weighted average of trajectory importance. Furthermore, we establish a novel worst-case lower bound for POSG policy optimization and offer a theoretical guarantee of performance improvement. Extensive experimental evaluations demonstrate that POSG surpasses other baseline algorithms across a discrete grid-world maze and three continuous locomotion control tasks.
@@
    POSG can use a single state-only demonstration trajectory to solve RL’s exploration problem and improve RL’s sample efficiency, demonstrated experimentally in Section VII-D1.
@@
The rest of this article is organized as follows. Section II describes some important progress of the recent related work. Section III briefly introduces the related preliminary knowledge. Then, the method we propose is introduced in detail in Section IV. Section V describes the theoretical analyses of POSG. The environmental settings are introduced in Section VI. In Section VII, we experimentally demonstrate the feasibility and effectiveness of the proposed algorithm in terms of exploration efficiency and learning speed. Finally, we summarize the main work of this study in Section VIII.
II Related work
@@
Learning from Demonstrations. LfD combines RL with expert demonstration data to improve policy exploration and accelerate learning. LfD trains a behavioral policy with a state-action visitation distribution similar to the demonstrator’s to accomplish this goal. Early LfD algorithms sample replay buffers containing expert demonstrations and self-generated data simultaneously to enhance the learning ability of the agent [24, 25]. Self-imitation learning (SIL) methods [33, 34, 27] train the agent to imitate its own past experiences only when the return of the previous episode is greater than the value estimate of the agent or returns of trajectories in the replay buffer. Episodic reinforcement learning methods utilize episodic memories to estimate the value of the state precisely and propagate its value to the previous states [35, 36, 37, 38, 39, 40]. Many other previous works introduce novel methods that enable the agent learns a range of diverse exploratory policies based on episodic memory [13, 41]. This study uses demonstration experiences in the state-action space to construct a shaping function, which can easily be integrated with existing RL methods, such as PPO [42].
@@
State-Only Imitation Learning (IL). Imitation learning aims to learn a control policy that imitates the behaviors of experts and outputs the same action when the agent receives the observation that occurred in the demonstration data set. The state-only IL method is a branch of imitation learning where the requirement for the demonstrator’s action information is alleviated. This approach expands the scope of the realistic application of IL. GAIfO [43] proposes to learn a policy to output actions that lead to similar effects as demonstrations. I2L [44] and SAIL [45] train the agent by minimizing the distance between state visitation distributions of the current policy and demonstrations. AILO [46] leverages demonstrations only containing observations to train an intermediary policy whose state transitions are close to the expert dataset. Our method can only utilize a few state-only demonstrations to achieve approximate credit assignments and avoid introducing extra neural networks.
@@
Metrics to Compute the Distance between Policies or Trajectories. Many studies propose to compute the distance between policies using Kullback–Leibler (KL), Bregman, or $f$-divergence [47, 48, 49]. Rényi divergence is introduced to compute the discrepancy of state visitation between different trajectories [50]; however, this divergence is a parametric distance measure, and this approach employs $k$-NN estimator to estimate the parameters of distributions efficiently. MADE maximizes the derivation of state-action visitations of the current policy from the explored regions of previous trajectories [51]. However, MADE relies on estimating the state-action visitation density, which can be non-trivial in high-dimensional control tasks. Similar to our method, CQL [52] and MCPO [14] adopt the MMD metric as the discrepancy measure; however, these algorithms require both state and action information of demonstrations to update the policy parameter.
@@
Credit Assignment. Various works have focused on the credit assignment problem, and credit assignment methods can be integrated with existing RL algorithms easily [53]. These methods assist us in understanding the association between sparse or delayed rewards and state or state-action pairs and reduce learning time by providing dense and supplemental rewards [54]. Zheng et al. [55] formulate pairwise weight functions of the state where the action is performed, the future highly-rewarded state, and the time horizon between the two states, which is learned by a special meta-gradient. Hindsight Credit Assignment (HCA) [18] converts the credit assignment problem into a supervised learning task by learning a hindsight probability function of actions. Counterfactual Credit Assignment (CCA) [19] uses a value function baseline with a hindsight information vector to implement credit assignment implicitly while avoiding giving away information about the agent’s actions to reduce potential bias. In RUDDER [22], an LSTM network is used to learn to redistribute the return of a trajectory to the preceding states before the final rewarding state. Episodic Backward Update [38] and Neural Episodic Control [35] enable efficient reward propagation by sampling from episodic memories and updating the value of all transitions more quickly. State Associative Learning [56] propagates credit directly by learning associations between states and arbitrary future states. Xu et al. [57] introduce an online meta-learning method to learn hyper-parameters of a discount $\gamma$ and bootstrapping parameter $\lambda$ of Temporal Difference (TD) for credit assignment. In contrast, our method does not incur high computational costs by dispensing with training auxiliary networks and can be regarded as a simple weighted return decomposition.
III Preliminaries
@@
This paper studies the credit assignment problem of reinforcement learning in tasks with sparse or delayed rewards. Before describing our method in detail, we introduce some preliminary knowledge about RL in this section. Then, the definition of Maximum Mean Discrepancy (MMD), which plays a vital role in our method, is provided.
III-A Reinforcement Learning
@@
III-B Maximum Mean Discrepancy
In this paper, we treat trajectories as deterministic policy distributions and use the maximum mean discrepancy (MMD) to empirically measure the difference (or similarity) between trajectories [58, 59, 60]. MMD is a non-parametric test statistic, and different from Kullback-Leibler (KL), Jensen–Shannon (JS), or $f$-divergences, calculating the MMD metric does not require the distribution parameters. This property makes it suitable for the problem considered in this study, where trajectories are treated as deterministic policies. The definition of MMD depends on the choice of function space.
Assume that the probability distributions $p$ and $q$ are defined on the space $\mathbb{X}$, and let $x$ and $y$ be the elements sampled from $p$ and $q$, respectively. Given a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ with the kernel function $k(\cdot,\cdot)$; we define the MMD as follows:
$\mathrm{MMD}^{2}(p,q,\mathcal{H})=\mathbb{E}[k(x,x^{\prime})]-2\mathbb{E}[k(x,y)]+\mathbb{E}[k(y,y^{\prime})],$ (5)
where $x,x^{\prime}\ \mathrm{i.i.d.}\sim p$ and $y,y^{\prime}\ \mathrm{i.i.d.}\sim q$. Using Eq. (5) to estimate the distance between $p$ and $q$ is tractable because of the inherent property of RKHS [58, 59].
III-C Trajectory-Space Smoothing
@@
IV Proposed Approach
This section presents smooth policy optimization (POSG), a computationally efficient framework to achieve approximate long-term credit assignments. The key insight is that we propose to exploit state-only demonstration trajectories to measure the relative importance of current state-action pairs. Our approach fuses the distributional information of demonstrations with the return information of trajectories.
IV-A State-Only Trajectory Distance
In this paper, we treat trajectories as deterministic policy distributions and use the maximum mean discrepancy to empirically measure the difference (or similarity) between trajectories [58, 59, 60]. Given a current trajectory $\tau$ and a state-only demonstration trajectory $\tau_{E}$, the nonparametric test statistic, the maximum mean discrepancy, can be used to measure the distance between them:
$\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})=$ $\underset{o,o^{\prime}\sim\rho_{\tau}}{\mathbb{E}}\left[k\left(o,o^{\prime}\right)\right]$ (8)
$-2\underset{\begin{subarray}{c}o\sim\rho_{\tau}\\ o_{E}\sim\rho_{E}\end{subarray}}{\mathbb{E}}\left[k(o,o_{E})\right]$
$+\underset{o,o_{E}^{\prime}\sim\rho_{E}}{\mathbb{E}}\left[k(o,o_{E}^{\prime})\right],$
where $k(\cdot,\cdot)$ is the kernel of a reproducing kernel Hilbert space $\mathcal{H}$, $\rho_{\tau}(\cdot)$ and $\rho_{E}(\cdot)$ are the state visitation distributions of $\tau$ and $\tau_{E}$, and $o,o^{\prime}$ and $o_{E}$, $o_{E}^{\prime}$ are observations sampled from $\rho_{\tau}$ and $\rho_{E}$, respectively. In practice, the function $k(\cdot,\cdot)$ in Eq. (8) is often defined as:
$k(x,y)=K\left(g(x),g(y)\right).$ (9)
Formally, the function $g$ gives us the feasibility that adapts the focus of the ${\rm MMD}$ metric for different aspects to different downstream tasks. In our experiments, we define the distance $D(\tau,\mathcal{M}_{E})$ from trajectory to the demonstration dataset $\mathcal{M}_{E}$ as follows:
This definition of $D(\tau,\mathcal{M}_{E})$ can only concern a relevant subset of the information in state observations. For example, we choose this information subset to be the coordinates $c=(x,y)$ of its center of mass (CoM) in the Key-Door-Treasure domain, i.e., the function $g$ maps a state observation $o$ to $c=(x,y)$.
IV-B Trajectory Importance Evaluation Mechanism
This section defines the trajectory importance based on a novel trajectory-level distance measurement and the trajectory return. In this manner, this importance integrates the state distribution information of the demonstrations with the reward signals of the trajectories. We can then fuse these two pieces of information into the guidance reward by a smooth weighted average of the trajectory importance. Specifically, we introduce a trajectory weight function $\omega_{\pi}(\tau|\mathcal{M}_{E})$ in Eq. (11) based on the MMD distance metric in the trajectory space and the conditional distribution $p_{\pi}(\tau|x)$. Mathematically, an exponential function $e^{-kd}$ is adopted by $\omega_{\pi}(\cdot|\mathcal{M}_{E})$ to smooth the weight of the trajectory, and $\omega_{\pi}(\cdot|\mathcal{M}_{E})$ is expressed as follows: